the formula for converting degrees fahrenheit(f) to degrees (c) is
c=5/9(f-32)Find c when f=68

Answer by oscargut(686)

c=5/9(f-32)then c when f=68 is 5/9 (68-32)=(5/9)(36)=20
Answer: 20 degrees (c)